### v2

# The early history of symmetric teleparallel gravity: An overlooked period

Abstract
It is noteworthy that symmetric teleparallel gravity has attracted considerable attention in recent years. A survey of the literature indicates that this surge of interest became particularly prominent around 2017 and 2018. However, together with my students and collaborators, we published a series of systematic and pioneering papers on this subject between 2004 and 2013. This article aims to clarify the early geometric development of symmetric teleparallel gravity preceding the recent surge of interest in $f(Q)$ gravity. For the sake of completeness and coherence, we will also briefly review our work on this topic carried out after 2018. In the final paragraph, we will write our personal perspectives on the future of symmetric teleparallel geometry.
Keywords: Curvature, torsion, non-metricity, metric-affine geometry, symmetric teleparallel gravity.
The author of this letter is solely myself; however, throughout the text he employs the pronoun “we” as the subject. Accordingly, the term “we” in this letter may sometimes refer only to me and, at other times, to us collectively in the literal sense. The reader will determine the intended referent from the context of the discussion.
From the earliest civilizations, the study of the heavens emerged from the practical needs of calendrics, navigation, and ritual timekeeping, gradually evolving into a quantitative science. Ancient Mesopotamian astronomers systematically recorded planetary motions and eclipses using numerical schemes, establishing the first predictive astronomy. Greek natural philosophers then transformed these empirical records into geometric cosmology: beginning with Eudoxus’ nested spheres and culminating in the geocentric synthesis of Claudius Ptolemy, celestial phenomena were modeled through combinations of uniform circular motions designed to preserve philosophical ideals of perfection in the heavens. During the Islamic Golden Age, astronomers critically refined Ptolemaic models through improved observations, trigonometric methods, and mathematical critiques that subtly separated physical reality from purely kinematic description. The Renaissance marked a decisive conceptual shift when Nicolaus Copernicus proposed a heliocentric ordering of the planetary system, later supported empirically by Tycho Brahe’s precision observations and mathematically reformulated by Johannes Kepler, who discovered that planetary orbits are ellipses governed by quantitative laws relating orbital periods and distances. These developments replaced qualitative cosmology with dynamical regularities and prepared the intellectual ground for a unified physics of the heavens and the Earth. By the late seventeenth century, this long evolution culminated in Isaac Newton’s synthesis, where celestial motion was no longer explained by geometric constructions alone but derived from universal dynamical principles acting through gravitation, thereby transforming astronomy into celestial mechanics.
When Isaac Newton published his theory of gravitation in 1687 in the Philosophiæ Naturalis Principia Mathematica, it successfully explained all astronomical observational data available at the time [1]. Subsequently, the Lagrangian [2] and Hamiltonian [3, 4] formulations of the theory were successfully constructed. Nevertheless, with the development of experimental techniques toward the end of the twentieth century, high–precision experiments concerning the measurement of the speed of light and the orbit of the planet Mercury placed Newton’s theory of gravitation under serious strain. On the other hand, Newtonian gravity was overshadowed by the symmetry structure underlying Maxwell’s electromagnetic theory [5, 6]. In short, the limitations of Newton’s gravitational theory, which had once appeared complete, had by then become unmistakably evident.
Einstein’s theory of gravitation, namely General Relativity (GR), then entered the stage [7, 8]. Naturally, it successfully overcame the observational and theoretical difficulties at which Newtonian gravitation had stalled. However, as experimental techniques continued to advance, GR proved insufficient to account for two major astrophysical observations. The first discrepancy appeared in the graphs of the linear velocities of stars located in the outer arms of spiral galaxies as a function of their distance from the galactic center. This shortcoming of GR is now known as the flatness problem of galactic rotation curves, or equivalently, the dark matter problem [9]. The second concerns the observed increase in the expansion rate of the Universe. This inadequacy of GR is presently referred to as the cosmic acceleration problem, or the dark energy problem [10]. On the other hand, a fundamental incompatibility between GR and quantum theory, established during the first quarter of the twentieth century, still remains unresolved [11]. In short, although Einstein’s theory of gravitation successfully resolved the problems of its time when it was first introduced, GR no longer stands as unchallenged as it once did.
It therefore follows naturally that GR, expressed by equating the Einstein tensor to the energy–momentum tensor, must be modified. If such a modification is implemented on the energy–momentum side of the field equations, one is typically led to postulate a specific form or distribution of matter tailored to resolve each individual observational problem. The dark matter and dark energy hypotheses fall into this category. Approaches of this kind generally provide only palliative remedies rather than a unified and fundamental solution.
An alternative strategy is to modify the geometric (left-hand) side of the Einstein field equations, an approach commonly referred to as modified theories of gravity. In some cases, the additional terms arising from such modifications are partially reabsorbed into an effective energy–momentum tensor, thereby altering its structure as well. A wide variety of alternative gravitational theories have been proposed in the literature, including $f(R)$ gravity, scalar–tensor theories, higher-curvature corrections, non-Riemannian extensions and metric–affine theories, massive gravity, MOND and its relativistic extensions, among others. For a comprehensive and up-to-date review, the reader may consult [12].
The rapid development of symmetric teleparallel gravity after 2017–2018 has created a substantial body of literature whose conceptual roots extend to earlier geometric formulations. A systematic reconstruction of this early stage is therefore useful for establishing continuity in the development of the subject and for clarifying the origin of several structures that reappear in modern formulations.
In this letter, we briefly summarize our pioneering contributions, carried out together with our collaborators between 2004 and 2013, from the perspective of the symmetric teleparallel theory of gravity, which constitutes a particular subclass within the broader framework of non-Riemannian theories of gravity or metric-affine gravity. Although a recent and extensive review article on symmetric teleparallel gravity exists [13], we noticed that our earlier works were entirely overlooked in that review. This observation motivated us to write the present letter.
The first work published on this subject is the short paper by James M. Nester and Hwei-Jang Yo, Symmetric teleparallel general relativity [14]. In that work, the authors attribute the gravitational force to the non-metricity sector of differential geometry with zero-curvature and zero-torsion. Essentially, they exploit the gauge freedom inherent in metric-affine geometry by setting the affine connection to zero in the coordinate basis. It is a kind of gauge fixing. Then, by making the non-Riemannian Einstein–Hilbert Lagrangian they could directly write down a Lagrangian for symmetric teleparallel general relativity that is equivalent to general relativity. However, they have not addressed or discussed the explicit form of their Lagrangian as a quadratic expression in the non-metricity tensor. They chose to formulate all computations entirely in tensor component notation.
In the paper by myself and Özcan Sert, A Solution to Symmetric Teleparallel Gravity [15], the STPG equations equivalent to Einstein’s equations were explicitly written down. To obtain solutions, a spherically symmetric static metric and, independently of the metric, a spherically symmetric static affine connection were proposed. The unknown functions appearing in these ansatzes were uniquely determined by the zero-curvature condition, the zero-torsion condition, and the Einstein-equivalent STPG field equation. In this way, a spherically symmetric static solution of the Einstein-equivalent STPG equations was obtained, and its singularity structure was analyzed. The name “symmetric teleparallel gravity” and the abbreviation “STPG” were used for the first time in this work. All calculations were carried out in the language of exterior algebra and within an orthonormal coframe formalism.
In the paper by myself, Mestan Kalay and Özcan Sert, Lagrange Formulation of the Symmetric Teleparallel Gravity [16], an STPG Lagrangian quadratic in the non-metricity and containing five independent coupling constants was presented for the first time. The variational STPG field equation was derived. Again, to find solutions, a spherically symmetric static metric and an independent spherically symmetric static affine connection were assumed, and the unknown functions were determined explicitly by the zero-curvature condition, the zero-torsion condition, and the STPG field equations, which in this case were not required to be Einstein-equivalent. It was shown that different choices of the free coupling constants lead to different families of solutions. All calculations were performed using exterior algebra in an orthonormal coframe.
In the paper by myself and Tekin Dereli, The Quadratic Symmetric Teleparallel Gravity in Two Dimensions [17], it was shown that while general relativity in (1+1) dimensions is trivial, STPG defines a genuinely dynamical theory, thereby highlighting an important advantage of STPG over GR. To demonstrate this, an STPG Lagrangian quadratic in the non-metricity with four coupling constants was constructed, the corresponding variational field equations were derived, and conformal and cosmological solutions were obtained. Moreover, it was shown that for special values of the four coupling constants, the STPG Lagrangian becomes equivalent to the Einstein–Hilbert Lagrangian in two dimensions. This allowed a clear distinction between GR-equivalent and non-GR-equivalent solution sectors. All calculations were again carried out in exterior algebra using an orthonormal coframe.
In the paper by myself, The Symmetric Teleparallel Gravity [18], it was shown for the first time that the five-parameter STPG Lagrangian previously introduced becomes equivalent to the Riemannian Einstein–Hilbert Lagrangian in (1+3) dimensions for specific values of the coupling constants. In addition, it was explicitly demonstrated that the affine connection can be expressed entirely in terms of metric components by exploiting the gauge freedom. This work can be regarded as the first to clearly establish symmetric teleparallel geometry as a metric geometry. Muzaffer Adak referred to this special choice as the natural gauge or inertial gauge; later, Tomi Koivisto and his collaborators introduced the term coincident gauge for the same construction[19]. Finally, by employing the natural gauge, families of spherically symmetric static solutions of the variational STPG field equations were obtained, not all of which are necessarily equivalent to GR. All calculations were performed in the framework of exterior algebra and orthonormal coframe.
In the paper by myself, Özcan Sert, Mestan Kalay and Murat Sarı, Symmetric Teleparallel Gravity: Some Exact Solutions and Spinor Couplings [20], the metric formulation of STPG—made possible by the natural gauge developed earlier—was systematically employed in the variational field equations derived from the quadratic non-metricity Lagrangian with five free coupling constants. Exact conformal, spherically symmetric static, cosmological, and pp-wave solutions were explicitly obtained. Such a broad class of exact solutions in STPG was presented for the first time in the literature. Once again, it was shown explicitly that for special values of the five coupling constants, the STPG Lagrangian is equivalent to the Einstein–Hilbert Lagrangian, allowing the solution families to be classified into GR-equivalent and non-GR-equivalent sectors. Finally, the coupling of spin-1/2 fermions to STPG was discussed for the first time in the literature by us. All calculations were carried out using exterior algebra in an orthonormal coframe formalism.
After this period, we suspended our research on STPG, since our earlier works received almost no attention within the alternative gravity community. Subsequently, beginning around 2017–2018, a series of influential papers on symmetric teleparallel gravity and later on $f(Q)$ gravity were produced by Tomi Koivisto, Alexey Golovnev, Jose Beltran Jimenez, Lavinia Heisenberg, and their collaborators. The recent resurgence of interest in symmetric teleparallel formulations, particularly through coincident general relativity and $f(Q)$ gravity, motivates a reexamination of earlier geometric approaches. Following these developments, the number of studies in the field of STPG increased rapidly [13, 19, 21, 22] and the references therein, as well as the works citing them.
We subsequently resumed our investigations in this field. In [23], we analyzed the gauge structure of the symmetric teleparallel theory of gravity. We also further developed our earlier work on the autoparallel curves of symmetric teleparallel geometry and, together with C. Pala, revisited in [24] the motion of a test particle within the STPG framework. In that study, we derived the corresponding trajectories and visited, in this context, the dark matter problem encountered in galactic dynamics.
In general, if non-metricity is present in a given geometry, the scalar product of two vectors becomes dependent on the path along which the vectors are parallel transported. For this reason, the physical viability of models constructed on non-metric geometries is sometimes called into question. In particular, within the framework of GR, this would imply that the proper time measured by a clock carried by an observer depends on the past trajectory of the clock. In other words, the measured time would depend on the history of the clock itself. This phenomenon is known as the second clock effect. However, experiments do not detect such an effect. We addressed the question of whether gravitational theories containing non-metricity can be freed from this apparent pathology and answered this question in the affirmative in our work [25], published together with our collaborators.
In order to reveal the dynamical degrees of freedom—one of the aspects in which the STPG theory exhibits advantages over GR—we revisited, together with our collaborators, the STPG framework in $(1+2)$ dimensions in [26]. During the same period, we also demonstrated for the first time, again in collaboration with our colleagues, that general teleparallel geometry admits a metric formulation obtained through a gauge-fixing procedure analogous to that encountered in symmetric teleparallel geometry [27]. While we had previously discussed minimal couplings of matter fields within the STPG framework, in [28] we investigated the non-minimal coupling between the STPG field and the Maxwell field. In our most recent work, we studied solar neutrino oscillations in an axially symmetric, weak, and slowly rotating STPG background and derived phenomenological upper bounds on four of the six free coupling constants appearing in the generalized Dirac equation [29].
Up to this point, we have summarized our contributions to the field of symmetric teleparallel gravity. More recently, we have been striving to apply the geometrical insight and computational techniques developed in this context to the theory of defects observed in crystalline structures. In the mainstream differential–geometric description of crystal defects, curvature is typically associated with disclinations, torsion with dislocations, and point defects with the trace of non-metricity. In contrast, we formulate all types of defects entirely within the framework of general teleparallel geometry [30, 31]. Using the model developed in this approach, we continue to work on projects addressing certain concrete problems studied in mechanical engineering, seeking symmetric teleparallel geometric solutions to them. We anticipate that symmetric teleparallel geometry will find increasing applications in physics, engineering, health sciences, and finance in the future, and we are actively pursuing research in these directions.
The purpose of this article is, therefore, not merely historical documentation but the clarification of the conceptual lineage of symmetric teleparallel geometry and its relation to ongoing research programs.
As a final remark, it is worth stating this point explicitly. While we carry out all calculations within the orthonormal coframe formalism using the language of exterior algebra, most researchers prefer to formulate their computations entirely in tensor component notation. It is therefore possible that the visibility of our work has not reached the level it deserves, partly due to the mathematical framework we have chosen to employ.
Acknowledgements: This work has been supported by TÜBİTAK (The Scientific and Technological Research Council of Türkiye) with the grant id 124F325.
Data Availability Statement: Data sharing not applicable to this article as no datasets were generated or analyzed during the current theoretical research.
Conflict of Interest: The authors declare no conflict of interest.